a unit vector in xy -plane that makes an angle of 45 degree with the vector i+j and an angle of 60 degree with the vector 3i-4j is

Grade:12

1 Answers

Prudhvi teja
13 years ago

Dear vikas

let vector b ai+bj

a.1+b.1 = 1*(2^0.5).cos45

= a+b =1

and a^2+b^2 = 1

=>(a+b)^2 - a^2-b^2 = 0

=>2ab = 0

so a=0 or b=0

so a=1,b=0

or a=0,b=1

a.3-b.4 = 5*cos60

=>3a-4b =5/2

either a=1 or b=1 doesnt suit.So there is no vector in xy-plane which satisfies given conditions

 

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.

All the best.

Regards,
Askiitians Experts
Prudhvi Teja

 

 


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian